# Parabolic Equations with Singular Coefficients and Boundary Data: Analysis and Numerical Simulations

### v1

@@ 4. Very weak solutions
$\qquad\leq\frac{C}{\alpha}\Bigg(\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\qquad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\qquad\Big(\|u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\tilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\tilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)\Bigg).$
Considering the negligibility of the nets $\tilde{a}_{\varepsilon}-a_{\varepsilon}$, $\tilde{b}_{\varepsilon}-b_{\varepsilon}$, $q_{\varepsilon}-\tilde{q}_{\varepsilon}$, $u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon}$, $g_{\varepsilon}-\tilde{g}_{\varepsilon}$, $f_{\varepsilon}-\tilde{f}_{\varepsilon}$, we get
$\qquad\leq\frac{C_{0}}{\alpha}\Bigg(\varepsilon^{-N_{0}}\Big(C_{1}\varepsilon^{-N_{1}}+C_{2}\varepsilon^{-N_{2}}+\frac{1}{\sqrt{\alpha}}\left(C_{3}\varepsilon^{-N_{3}}+C_{4}\varepsilon^{-N_{4}}+C_{5}\varepsilon^{-N_{5}}\right)\Big)$
$\qquad\times\Big(1+\frac{1}{\sqrt{\alpha}}\left(C_{6}\varepsilon^{-N_{6}}+C_{7}\varepsilon^{-N_{7}}\right)+C_{8}\varepsilon^{-N_{8}}\Big)\big(C_{9}\varepsilon^{N_{9}}+C_{10}\varepsilon^{N_{10}}\varepsilon^{-N_{11}}$
$\qquad+C_{11}\varepsilon^{N_{12}}\varepsilon^{-N_{13}}+C_{12}\varepsilon^{N_{14}}\varepsilon^{-N_{15}}+C_{13}\varepsilon^{N_{16}}\varepsilon^{-N_{17}}+C_{14}\varepsilon^{N_{18}}+C_{15}\varepsilon^{N_{19}}+C_{16}\varepsilon^{N_{20}}\big)\Bigg)$
for some $C_{j}>0$, j=0,,16j=0,...,16, $N_{k}\in\mathbb{N}$, $k=0,\ldots,8,\,11,\,13,\,15,\,17$, and all
$N_{9},\,N_{10},\,N_{12},\,N_{14},\,N_{16},\,N_{18},\,N_{19},\,N_{20}\in\mathbb{N}$. Then, for some $C_{D}>0$ and all $K\in\mathbb{N}$
$\|U_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\leq C_{D}\varepsilon^{K},$
and this completes the proof of Theorem 4.2. ∎
@@
$b\in L^{\infty}([0,T];W^{1,\infty}(0,1))$, $q\in L^{\infty}(0,1)$, $g_{0},\,g_{1}\in H^{1}([0,T])$, and let $a_{\varepsilon}$ be any $W^{1,\infty}$-regularisation of $a$, $b_{\varepsilon}$ be any log-$L^{\infty}(W^{1,\infty})$-regularisation of $b$, $(q_{\varepsilon})$ be any $L^{\infty}$-regularisation of $q$, $g_{0,\varepsilon}$, $g_{1,\varepsilon}$ are any time-distributional regularisations of $g_{0}$ and $g_{1}$ respectively, that is
@@
$\|q_{\varepsilon}-q\|_{L^{\infty}(0,1)}\to 0,\quad\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])}\to 0,\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])}\to 0$
as $\varepsilon\to 0$. Let the initial condition satisfies $u_{0}\in H^{1}(0,1)$. Let $u$ be a very weak solution of the initial/boundary problem (4.2). Then for any families $a_{\varepsilon}$, $b_{\varepsilon}$, $q_{\varepsilon}$, $g_{0,\varepsilon}$, $g_{1,\varepsilon}$, $u_{0,\varepsilon}$ and $f_{\varepsilon}$ such that
@@
as $\varepsilon\to 0$, any representative $(u_{\varepsilon})$ of the very weak solution converges as
$\|u(t,\cdot)-u_{\varepsilon}(t,\cdot)\|_{H^{1}([0,T];L^{2}(0,1))}+\|u(t,\cdot)-u_{\varepsilon}(t,\cdot)\|_{L^{2}([0,T];H^{2}(0,1))}\to 0$
for $\varepsilon\to 0$ to the unique classical solution $u\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1))$ of the initial/boundary problem (4.2) given by Theorem 3.1.
Proof.
Let us denote $\tilde{U}_{\varepsilon}$ as the difference between $u$ and $u_{\varepsilon}$, that is
$\tilde{U}_{\varepsilon}(t,x):=u(t,x)-u_{\varepsilon}(t,x).$
Then the net $\tilde{U}_{\varepsilon}$ is a solution to the initial/boundary problem
$\left\{\begin{array}[]{l}\partial_{t}\tilde{U}_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}\tilde{U}_{\varepsilon})(t,x)+b_{\varepsilon}\partial_{x}\tilde{U}_{\varepsilon}+q_{\varepsilon}(x)\tilde{U}_{\varepsilon}(t,x)=F_{\varepsilon}(t,x),\\ (t,x)\in[0,T]\times(0,1),\\ \tilde{U}_{\varepsilon}(0,x)=(u_{0}-u_{0,\varepsilon})(x),\quad x\in(0,1),\\ \tilde{U}_{\varepsilon}(t,0)=(g_{0}-g_{0,\varepsilon})(t),\quad\tilde{U}_{\varepsilon}(t,1)=(g_{1}-g_{1,\varepsilon})(t),\quad t\in[0,T],\end{array}\right$. (4.9)
where
$+\big(q_{\varepsilon}(x)-q(x)\big)u(t,x)+f(t,x)-f_{\varepsilon}(t,x)$
Analogously to Theorem 4.2 we have that
$\|\tilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\tilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
$\qquad\leq\frac{C}{\alpha}\Bigg(\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\qquad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\qquad\times\Big(\|u_{0}-{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|{a}_{\varepsilon}-a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{x}({a}_{\varepsilon}-a)\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|{b}_{\varepsilon}-b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|{q}_{\varepsilon}-q\|_{L^{\infty}(0,1)}\|{u}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f-{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|g_{0}-{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1}-{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)\Bigg).$
Since
$\|g_{0}-g_{0,\varepsilon}\|_{H^{1}([0,T])},\quad\|g_{1}-g_{1,\varepsilon}\|_{H^{1}([0,T])},\quad\|f-f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\to 0$
for $\varepsilon\to 0$ and $u$ is a very weak solution of the initial/boundary problem (4.2), we get
$\|\tilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\tilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\to 0$
for $\varepsilon\to 0$. This proves Theorem 4.3.

### v2

@@ 3. Very weak solutions for the non-homogeneous Dirichlet conditions
$\quad\leq\frac{C}{\alpha}\Big(\exp{\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\right)\right\}}\Big(1+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\qquad+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
$\qquad\times\Big(\|u_{0,\varepsilon}-\widetilde{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|\widetilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{x}(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\widetilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\widetilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\widetilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|g_{0,\varepsilon}-\widetilde{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}-\widetilde{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)$
$\qquad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0,\varepsilon}-\widetilde{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}-\widetilde{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\big)\Big).$
By Proposition 3.2 (applied to $\widetilde{u}_{\varepsilon}$) and the moderateness assumptions, the nets
$\|\widetilde{u}_{\varepsilon}\|_{L^{2}(0,T;L^{2})},\qquad\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}(0,T;L^{2})},\qquad\|\partial_{x}^{2}\widetilde{u}_{\varepsilon}\|_{L^{2}(0,T;L^{2})}$
are moderate. On the other hand, by the negligibility assumptions in Definition 8, the nets
$\|\widetilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}W^{1,\infty}},\quad\|\widetilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}W^{1,\infty}},\quad\|\widetilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}},\quad\|f_{\varepsilon}-\widetilde{f}_{\varepsilon}\|_{L^{2}(0,T;L^{2})}$
are negligible. Hence, each term on the right-hand side of (3) is negligible, and therefore
$\|F_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}\ \text{is negligible.}$
Similarly, $U_{\varepsilon}(0,\cdot)=u_{0,\varepsilon}-\widetilde{u}_{0,\varepsilon}$ is $H^{1}$-negligible and the boundary differences $U_{\varepsilon}(\cdot,0)$ and $U_{\varepsilon}(\cdot,1)$ are $H^{1}(0,T)$-negligible.
Applying estimate (3.2) to (3.12) yields that
$\|U_{\varepsilon}\|_{H^{1}(0,T;L^{2}(0,1))}+\|U_{\varepsilon}\|_{L^{2}(0,T;H^{2}(0,1))}$
is negligible. This proves uniqueness of the very weak solution in the sense of Definition 8. ∎
@@
$b\in L^{\infty}([0,T];W^{1,\infty}(0,1))$, $q\in L^{\infty}(0,1)$, $g_{0},\,g_{1}\in H^{1}([0,T])$. Assume moreover that $u_{0}\in H^{1}_{0}(0,1)$. Let $u$ be the unique classical (weak) solution of (3.1), i.e.
$u\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1)).$
Then for any families $a_{\varepsilon}$, $b_{\varepsilon}$, $q_{\varepsilon}$, $g_{0,\varepsilon}$, $g_{1,\varepsilon}$, $u_{0,\varepsilon}$ and $f_{\varepsilon}$ such that
@@
$\|q-q_{\varepsilon}\|_{L^{\infty}(0,1)}\to 0,\quad\|u_{0}-u_{0,\varepsilon}\|_{H^{1}_{0}(0,1)}\to 0,$
as $\varepsilon\to 0$, any representative $(u_{\varepsilon})$ of the very weak solution converges to $u$ in the sense that
$\|u-u_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|u-u_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\to 0,\qquad\varepsilon\to 0.$
Proof.
Let us denote $\widetilde{U}_{\varepsilon}$ as the difference between $u$ and $u_{\varepsilon}$, that is
$\widetilde{U}_{\varepsilon}(t,x):=u(t,x)-u_{\varepsilon}(t,x).$
Then the net $\widetilde{U}_{\varepsilon}$ is a solution to the initial/boundary problem
$\left\{\begin{array}[]{l}\partial_{t}\widetilde{U}_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}\widetilde{U}_{\varepsilon})(t,x)+b_{\varepsilon}\partial_{x}\widetilde{U}_{\varepsilon}+q_{\varepsilon}(x)\widetilde{U}_{\varepsilon}(t,x)=F_{\varepsilon}(t,x),\\ (t,x)\in[0,T]\times(0,1),\\ \widetilde{U}_{\varepsilon}(0,x)=(u_{0}-u_{0,\varepsilon})(x),\quad x\in(0,1),\\ \widetilde{U}_{\varepsilon}(t,0)=(g_{0}-g_{0,\varepsilon})(t),\quad\widetilde{U}_{\varepsilon}(t,1)=(g_{1}-g_{1,\varepsilon})(t),\quad t\in[0,T],\end{array}\right$. (3.14)
where
$+\big(q_{\varepsilon}(x)-q(x)\big)u(t,x)+f(t,x)-f_{\varepsilon}(t,x).$
Since $u\in L^{2}([0,T];H^{2}(0,1))$, each term in $F_{\varepsilon}$ tends to zero in $L^{2}([0,T];L^{2}(0,1))$, hence $\|F_{\varepsilon}\|_{L^{2}}\to 0$.
Analogously to Theorem 3.4 we have that
$\|\widetilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\widetilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}$
$\quad\leq\frac{C}{\alpha}\Big(\exp{\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\right)\right\}}\Big(1+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\qquad+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
$\qquad\times\Big(\|u_{0}-{u}_{0,\varepsilon}\|_{L^{2}(0,1)}+\|a-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|\partial_{x}(a-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|b-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}{u}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|q-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|{u}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f-{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|g_{0}-{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1}-{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)$
$\qquad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0}-{g}_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1}-{g}_{1,\varepsilon}\|_{H^{1}([0,T])}\big)\Big).$
Since
@@
for $\varepsilon\to 0$ and $u\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1))$, we obtain
$\|\widetilde{U}_{\varepsilon}\|_{H^{1}([0,T];L^{2}(0,1))}+\|\widetilde{U}_{\varepsilon}\|_{L^{2}([0,T];H^{2}(0,1))}\to 0,$
which proves Theorem 3.5. ∎